Run Coordination Meetings:Spring2019 Run

From GlueXWiki
Jump to: navigation, search

Spring 2019 GlueX Run

Meeting General Agenda

Run Planning Meeting Notes, Feb. 6-Fev 13, 2019

Run Planning Meeting Notes, Feb. 14-Feb 21, 2019

Run Planning Meeting Notes, Feb. 21-Feb 27, 2019

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1] https://bluejeans.com/660743227.

2.) To join via Polycom room system go to the IP Address: 199.48.152.152 (bjn.vc) and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at 199.48.152.152
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Beam and Hall configuration

Information in blue is specifically for the DIRC commissioning

Information in magenta is specifically for the PrimEx-η run

  • Beam energy expected:
    • 11.630 GeV (10 MeV less than value for Fall 2016 to Fall 2018 up Nov. 15th, due to imbalanced linac setting).
  • Solenoid:
    • DIRC commissioning: 1350A
    • PrimEx-η run: off
  • CW Beam:
    • 250 MHz frequency.
    • Range:
      • DIRC commissioning: 50 nA-1.2 μA.
      • PrimEx-η run: 1 nA-500 nA.
    • Chopper slit: shared with Hall C until 03/10/2019. Then alone on slit A.
    • Standard current and radiators
      • DIRC: Standard commissioning current: 225 nA on 1.4·10-4 R.L. radiator (JD70-104 17 μm diamond).
      • DIRC: Hight intensity tests: 5 days at 1.2 μA on 1.4·10-4 R.L. radiator (JD70-104 17 μm diamond) (or 0.5 μA on JD70-105, 47 μm diamond of 3.9·10-4 R.L.)
      • PrimEx-η run: Standard production on 4He: 320 nA. on 10 µm Al (11.2·10-5 R.L.)
      • PrimEx-η run: Standard production on Be: 100 nA. on 10 µm Al (11.2·10-5 R.L.)
    • RSAD limit for most of the run, excluding high intensity tests: 1.3 μA on ·10-4 R.L. radiator for 5 days.
  • Position Slow Lock on or FFB on (FFB doesn't operate below 50nA).
  • Radiators:
    • Goniometer radiators:
      • Diamond JD70-104 (17 μm - 1.4·10-4 R.L., 7×7 mm²) - wider rocking curve - Tested in Spring 2018 and used at the beginning of Fall 2018. Probably has some radiation degradation.
      • Diamond JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) - Used for the second part of GlueX's Fall 2018 run. Good performance.
    • 40 μm Al radiator (44.9·10-5 R.L.)
    • V-wire (for TAC run).
    • Amorphous
      • 1.64 µm Al (1.86·10-5 R.L., see study)</font>
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Tagger quadrupole on (-4.2 A).
  • Collimator hole: Both 5 mm diameter and 3.4 mm diameter. Default running: 5.0 mm.
  • Miscellaneous:
    • CPP (Charged pion polarizability) shielding will be temporary installed.
    • Transition Radiation Detector in the Hall (Parasitic runs)

General Information

This document describes the run plan for the Hall D Spring 2019 run.

  • Schedule for the run (Updated 01/28/2019):
  1. Jan. 22nd - Feb 7th: Electron beam restoration. (Now pushed to Feb. 7th)
  2. Feb. 7thth - Feb. 20th: 2-weeks DIRC Commissioning (Solenoid at nominal field. LH2 target))
  3. Feb. 20th (tentatively!): 1 day of CPP test (Solenoid at nominal field. LH2 target)
    1. Move the FCAL platform downstream. Install the CPP shielding mount.
    2. Take data in 3 configurations: a) no shilding, b) 1 layer of shielding, 2 layers.
  4. Feb. 20th morning (tentatively!) - Apr. 15th: PrimEx-eta run . Tentative plan:
    1. Turn off the solenoid.
    2. Install the FCAL platform in the regular position. Remove the LH2 cell (leave only the shell). Remove ST. Reinstall the beam line.
    3. Take data for CCAL calibration at a low beam current. 3 shifts.
    4. Take PRIMEX data with no target. 1 shift.
    5. Intall Be target. Install ST (not connecting the cables). Fiducialize the Be target. Survey the target (is it needed?). Survey the CCAL in the central position.
    6. Take PRIMEX data. 5 shifts.
    7. Remove Be target, Install the LHe cell, install and connect the ST. Fiducialize the cell and survey the target.
    8. Take PRIMEX data with the empty LHe target (not cooling the target). 2 shifts.
    9. Cool and fill the LHe target.
    10. Take PRIMEX data till the end.


  1. Nominally, every Mondays: 12h of beam study. Every Tuesdays, if necessary: 8h of RF recovery.
  • Accelerator overall plan and priorities:
  1. Deliver physics beam to 4 halls. (Hall A stop early, on March. 10th. Halls B and or C will always be at 5th pass, so Hall D beam will always be RF-separated.)
  • DIRC commissioning page
  • PrimeX run plan
  • DAQ and Trigger high-intensity tests during DIRC commissioning.
  • CPP tests during last day of DIRC commissioning.
  • Parasitic TRD runs.
  • If opportunity arises, test fast raster beam envelope measurement with larger x-rastering (Richard Jones)
  • Strait track runs not expected to be needed

Runplan

  1. Verify electron beam quality and establish photon beam (assume 1 day).
    • Ion Chambers trip threshold must be set using Al. radiators
    • Tune electron beam with radiator retracted
      • Radiator (amorphous and diamond) must be retracted. Collimator should be in blocking position.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 < 0.1 mrad/hr (tagger area, gammas, on the ground below the goniometer).
          • RAD102_P3 ~ 0.3 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 ~ 0. mrad/hr (Hall. Neutrons).
    • Establish good photon beam
    • Make sure tagger quadrupole is on and at -4.2 A. (It can be turned on only once the Hall is in power/beam permit.)
      • Insert 2*10-5 RL radiator. Call MCC and ask to mask the FSD and turn off beam each time a radiator is moved
        • Typical levels are (for 2×10-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 35 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= < 0.5 mrad/hr (tagger area, gammas, on the ground below the goniometer)
          • RAD102_P3= ~2 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Verify beam position and envelope (Hovanes).
      • Ask MCC to set beam position on 5C11b to x=-0.5 mm and y=+1.0 mm.
      • Ask MCC for 5C11 and 5C11B HARP. Then do our radiator HARP scan, see instruction. If it looks good, proceed to next step. Otherwise, inform MCC.
      • Insert profiler (no need for masking FSD when an object is inserted in the photon line), run with 10×10-4 radiator and 50 nA. Beam on the profiler should be at x=0.0±0.5cm; y=0.0±0.5cm. If not, inform MCC.
    • Radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 5 min with the 10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 3 min with the 3*10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
    • Photon flux optimization and Active Collimator calibration. (20 + 20 mins). Call Hovanes.
      • re-Insert the 10-4 radiator (Call MCC). Use about 100 nA beam current. This will provide the Act. Col. nominal positions for its high gain values.
      • Determine maximum photon beam transmission by doing 2D scans using the collimator x/y motions (Hovanes. 1h)
      • Repeat procedure with 2*10-5 radiator (Call MCC) and 50 nA beam current. This will provide the Act. Col. nominal positions for its low gain values.
      • Record on the white board what is the beam position on 5C11b and Active Collimator for optimal photon transmission, both for low and high Act. Col. gain.
      • Until Hovanes determines to optimal 5C11b and Active Collimator position, keep profiler in and beam locked (MCC's duty) on the nominal 5C11b and x=0,y=0 profiler position. Once Active Collimator nominal beam position is known, retract profiler and lock the beam (MCC) on the active collimator.
  2. Detector and beamline checkout.
    • PS detector threshold scan. (A. Somov, 1h) Can be done early if necessary: nominal beam quality not mandatory.
    • Microscope yield study (row-by-row scan). (Richard Jones/Jim McIntyre, 1h) Can be done early if necessary: nominal beam quality not mandatory.
      • Call Richard Jones 20 min before the start of this task. Call Alex Somov to set-up the PS trigger.
      • Insert 5mm collimator.
      • Procedure to follow
      • Any type of beam quality will do.
      • 10-4 radiator, 10-3 convertor, 80 nA beam current.
    • Reestablish L1 trigger Alex. Somov. 2h).
    • Reestablish DAQ (Sergei, 2h).
    • General detector calibration (Sean Dobbs. 1h). Use 10-4 radiator.
      • Run 15 min (solid time) in Long Mode (Mode 10)
      • Run 15 min (solid time) in Production Mode (Mode 9).
  3. Check 17 μm diamond (JD70-104) alignement. (We expect to use 2 polarization configuration only) (Hovanes. 1h).
  4. DIRC commissioning
    • DIRC commissioning page
    • Standard configuration: 17 μm diamond, 225 nA (1.2 μA for high intensity run), 5 mm hole, TPol on (for TPol syst. studies. Not necessary for DIRC) 75 μm TPol convertor??. 2h runs and no more, as it may be hard to correct for calibration drifts. Cycle over 3 sets of runs: ∣∣, then ⊥, then amorphous runs (i.e. two diamond settings only). DAQ start run overhead: 3min.
    • Luminosity scans on amorphous radiator. 50-300 nA. Use 10 µm Al (1.12·10-4 R.L.), on amorphous radiator ladder.
    • Tagger quadrupole on
    • Harp scan (30 min): Only once in the middle of the DIRC commissioning (beside the ones done during beam set-up).
    • One Empty target run (2h: 1- h to empty/fill back the target, 1h running). Standard production current.
    • Long mode 10 min run. Twice or Thrice a week. Beam must be off. On opportunistic basis if possible (Purpose: monitor neutron radiation damage to ST due to nuclear target higher neutron yield.)
    • No TAC run necessary
  5. Transition to PrimEx-η
    • CPP (Charged pion polarizability) tests (Not firm yet.)
      • To be set-up and done during the last day of DIRC allocated commissioning time.
      • Use LH2 target, keep solenoid on at nominal field, DIRC stays in nominal position.
      • Move downstream platform downstream
      • Install passive shielding in front of TOF.
      • Set-up and check specific CPP trigger: 2h (Alex Somov).
      • Tests (to be done for various thicknesses of passive shielding).
      • End of test: ramp down solenoid, remove shielding.
  6. PrimEx-η
    • PrimEx-η runplan page
    • Standard configuration: 10 µm Al (11.2·10-5 R.L. Beam current depends on target. 2h runs and no more, as it may be hard to correct for calibration drifts</font>. DAQ start run overhead: 3min.
    • Two TAC runs (to be done with the TAC detector, not with ComCal).
      • First one: shortly after March 10th, once we Hall A is down and we run on the Hall A slit.
      • Second one: near the end of the run.
      • HARP scan (only Hall D harp): 10 min. Every two days.
    • Long mode 10 min run. Twice or Thrice a week. Beam must be off. On opportunistic basis if possible (Purpose: monitor neutron radiation damage to ST due to nuclear target higher neutron yield.)

Hall D situation room (current run plan)

Hall D counting room white board


Runplan time charts

Spring 2019 time chart:

Spring19 week1.jpg

Spring 2019 commissioning/run time chart. Week 1.
The horizontal scale is the time during which beam is available (in other words, it would be the real time with 100% of beam efficiency).
The task time estimates given on the plot have 100% contingency.

FSpring19 week2.jpg

Run time chart. Week 2.

FSpring19 week3.jpg

Run time chart. Week 3.


Accelerator status

Hot Checkout

  1. Hot Check OUT Status 99% ready ; 1% checked ; 0% not ready


Readiness for Spring 19 run

  • Solenoid (Eugene/Beni)
    • Started HCO. Done
    • Will test ramp to 300 A the magnet. Done
    • Will ramp the magnet for the run once beam is close to delivery.
  • Surveys (Tim/Mark)
    • (ComCal will be surveyed during the DIRC-PrimEx transition)
    • (The Be target will be surveyed during the DIRC-PrimEx transition)
    • DIRC, was surveyed Wednesday, Jan 30th and checked again on Monday (02/04).
  • DAQ (Sergei).
    • Implementing DIRC's scalers and their transfer to EPICS. Done.
    • Beni started cosmic runs ⇒ Needed to power the detectors. Started on Jan. 30th. Things lookin good.
  • Trigger (Alex S.)
    • During DIRC commissioning, trigger will be the same as for GlueX production runs.
    • A LED trigger for DIRC need has been implemented and tested. It seems to be working. Justin checked that everything is fine.
    • DIRC readout has been integrated to the main physics trigger.
    • During PrimEx-Eta run, trigger will be the same as for ComCal Fall 2018 commissioning.
    • Implementing a new feature in SSP, to read the total energy from the crates for every 4ns bin. This would be for raw mode. The implementation is going well and may be available for this run.
    • Need to set-up the CPP trigger. This will be done during the day allocated to CPP tests. Testing the special trigger should take about 2h of beam time.
  • Beam line
    • Beam pipe
      • Old beam pipe is installed and will be used for DIRC commissioning
      • New beam pipe to be installed during transition DIRC-PrimEx transition, and used for PrimEx-Eta run. The old pipe will be removed during the transition and there will be no beam pine during the PrimEx run
    • Diamond status (Tim/Hovanes).
      • Discussion to install a low quality thick diamond (U.Conn. inventory) for DIRC commissioning: Schedule is too tight.
    • Tagger hodoscope (A. Somov)
      • Ready. There are several counters that are broken/have problems (see logbook entry 3643240. There is no spare nor time to fix them during the run. They will be dealt with at the calibration stage.
      • Sasha reminded us that the low-energy counters will need to be switched off during the DIRC high intensity runs.
    • Tagger microscope (Richard Jones/Beni)
      • Dark rate data taken Monday (Jan. 21st)
    • Goniometer (Hovanes).
      • Ready
    • Collimator Motion (Hovanes).
    • Active collimator detectors(Richard Jones/Alexandre Deur)
      • Act. Col. baselines have been zeroed mid July 2018 and checked on 08/15. Stable enough. They don't need to be revisited for the Spring 2019 run.
      • Ready
    • Profiler
    • luminosity monitor (Hovanes).
    • Halo monitors
      • They have not been touched. They should be ready. To be checked.
      • Should be ready
    • luminosity monitor
    • PS system (Alex Somov)
      • Ready.
    • Beam line magnets (PS magnet sweeping magnet and tagger quads).
    • Polarimeter (Sebastian Cole/M. Dugger).
    • LH2 Target (Tim/Alexandre D.)
      • Target has been cool-down tested. It is working fine.
    • Be target
    • 4He target
    • TAC (Hovanes/Alex Somov)
      • Plan to cured after the Fall 2018 run was not performed.
    • Slow control (Hovanes).
      • Working on DIRC. Missing Temperature voltages and callers from CODA. Some alarms are still missing.
      • ComCal. No new work needed. It is ready.
    • RF (Sergei)
      • Done: Checked with cosmic trigger. Looks fine
    • Radiation monitors (Alexandre D.)
      • Operational
  • Detector systems
    • ST (Mark Ito)
      • Ready and tested (Beni)
      • Will be used for DIRC commissioning. Will remain in place for PrimEx-Eta run and used for the 4He run, but will be turned off for the Be part of the run.
      • Random trigger has been implemented (for Beni to monitor neutron radiation damage).
    • FDC (Lubomir)
      • Got cosmic trigger to get data. Things look fine.
      • Gas chromatograph installed near the gas panel. Connected to the FDC and CDD output line.
      • Chiller kept tripping. The problem was fixed by disabling an interlock. The other problem (communication. Maybe due to a IOC) went away. Reason unknown. In all, the chiller is operational.
      • Lights for the camera near the bubbler/rotameter have been set-up (Hall's lights are dimmed because of DIRC).
    • CDC (Beni)
      • Up and running.
      • Flow controller replaced by a rotameter. It will be used from now on instead of the flow controller , unless we understand the origin of the green thing that clogged the flow controller.
    • BCal (Elton/Mark D.)
      • Did LED run on March 29th. It looks good.
      • Will start HCO
      • Low voltages issue (could not turn on) fixed: Two channels had the wrong parameter settings.
    • FCal (Collin Gleason)
      • BCal turned on fine. Need to turn on LED pulser to get some data.
      • One FADC module is dead and two modules have a bad channel. Sasha to replace the bad module on 02/06.
      • HCO done
    • ComCal (Alex Somov)
      • Taking test LED data. They look fine.
    • TOF (Mark Ito/Beni)
      • Ready: It was turned on and looked fine (Beni).
      • Checked baselines with cosmic data.
      • One HV broken was fixed (02/05) by Beni.
    • DIRC (J. Stevens)
      • Fixed a few issues in the box (01/20th).
      • Since 01/24th, everything is connected and readout with low rate trigger looks good.
      • Took high rate (LED) triggers. Looking at it. Need to keep taking LED data to set-up threshold for each channel. LED data will be taken until beam comes.
  • RSAD and ESAD Safety documents (Lubomir):
    • Documents approved and ready for shift takers to sign.
    • DIRC replied ro readiness review recommendations on 01/25th.
    • New ComCal OSP submitted and signed (the previous version expired March 2nd.)
  • Phones
    • System experts should have found their on-call phones by now.
    • System experts should have their phone charged.
    • System experts should start carrying their (charged) phone.
  • Hall Status/safety (Mark/Tim)
  • Other topics?